# Lissajous Varieties

### v1

@@ 1 Introduction
Let $A=\begin{pmatrix}a_{1}&a_{2}&\cdots&a_{n}\end{pmatrix}\in\mathbb{Z}^{d\times n}$ be such that ${\rm rank}(A)=d$ and $\mathbb{Z}A=\mathbb{Z}^{d}$. Let $P_{A}\subset\mathbb{R}^{d}$ be the polytope obtained as the convex hull of the lattice points $\{\pm a_{1},\ldots,\pm a_{n}\}\subset\mathbb{Z}^{d}$. The affine variety ${\cal L}_{A,b}$ is irreducible of dimension $d$. Moreover, for generic $b\in\mathbb{C}^{n}$, the degree of ${\cal L}_{A,b}$ is $\deg({\cal L}_{A,b})=2^{-{\rm CL}_{A}}\,d!\,{\rm vol}(P_{A})$, where ${\rm vol}(\cdot)$ denotes the euclidean volume and ${\rm CL}_{A}$ is the number of zero entries of a generic vector in the kernel of $A:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$.
The quantity $d!\,{\rm vol}(P_{A})$ is called the normalized volume of $P_{A}$. The polytope $P_{A}$ in Example 1.2 is the line segment [1,1][-1,1], with normalized volume two, which is the degree of the circle. The polygon $P_{A}$ in Example 1.3 is a hexagon with normalized volume $6=\deg{\cal S}_{A}$. The notion of “generic bb” in Theorem 1.4 will be made more precise in Section 2, and we give a formula for $\deg({\cal L}_{A,b})$ which holds for any $A,b$, but requires more notation (Theorem 2.5). We show that the hypersurface ${\cal C}_{A}$ obtained from the incidence matrix $A$ of the $n$-cycle graph $C_{n}$ is the cycle polynomial from [23, Section 4.2]. Using Theorem 2.5 and a result from [5], we prove a formula for the degree of the cycle polynomial (Proposition 2.10).
@@ 2 Dimension and degree
It remains to show that $S_{v}$ consists of $d!\,{\rm vol}(P_{A})$ points. This is an application of Kushnirenko’s theorem [12, Théorème III’], which predicts the maximal (and expected) number of solutions to a system of Laurent polynomial equations with identical monomial support. In order to apply this theorem, we must ensure that our equations are non-degenerate with respect to the polytope $P_{A}$. This means that none of the facial subsystems have a solution in $(\mathbb{C}^{*})^{d}$. By our assumption that ${\rm rank}(A)=d$, the centrally symmetric polytope $P_{A}$ is full-dimensional and it contains the origin in its interior. This implies that no face of $P_{A}$, except $P_{A}$ itself, contains both the lattice points $a_{j}$ and $a_{j}$. Therefore, there are no dependencies among the coefficients appearing in any facial subsystem. Hence, for generic $c_{jk}$, our equations are indeed non-degenerate and Kushnirenko’s upper bound $|S_{v}|\leq d!\,{\rm vol}(P_{A})$ is attained.
@@
We end the section by applying our degree formula to a family of Chebyshev hypersurfaces which arises when studying elliptopes of cycle graphs [21, 23]. This is relevant in semidefinite completion problems, see for instance [14]. Rephrasing [23, Equation (28)] in our notation, the $n$-th cycle polynomial $\Gamma_{n}^{\prime}$ is the defining equation of the Chebyshev hypersurface ${\cal C}_{A_{n}}$, with
@@
Notice that $\Gamma_{n}^{\prime}$ is defined up to scale. Its degree was computed up to $n=11$ in [23, Table 1]. In the paragraph preceding Conjecture 4.9 in [23], it is stated that a closed formula for $\deg(\Gamma_{n}^{\prime})$ is not known. The next proposition provides such a formula.
@@ 3 Determinantal equations
Fix $A\in\mathbb{Z}^{d\times n}$ and $\beta\in(\mathbb{C}^{*})^{n}$. Let $Y_{A,\beta}\subseteq\mathbb{C}^{n}$ be the corresponding scaled affine toric variety. If $g_{1},\dotsc,g_{r}$ generate the ideal of $Y_{A,\beta}\cap(\mathbb{C}^{*})^{n}$ in $\mathbb{C}[y_{1}^{\pm 1},\ldots,y_{n}^{\pm 1}]$, then
@@
It is easier to obtain generators for the ideal of $Y_{A,\beta}\cap(\mathbb{C}^{*})^{n}$ than for the toric ideal of $Y_{A,\beta}$. First, one computes a $\mathbb{Z}$-kernel of $A$. If this kernel is generated by $m_{1},\ldots,m_{r}\in\mathbb{Z}^{n}$, with $r=n-{\rm rank}(A)$, then the binomials $g_{1},\ldots,g_{r}$ in Theorem 3.1 are given by $y^{m_{k}}-\beta^{m_{k}},k=1,\ldots,r$ [24]. To avoid computing matrix inverses, one may clear denominators by writing $m_{k}=u_{k}-w_{k}$ with $u_{k},w_{k}\in\mathbb{N}^{n}$ and use $g_{k}=\beta^{w_{k}}y^{u_{k}}-\beta^{u_{k}}y^{w_{k}}$ instead.
@@
To prove Theorem 3.1, we first state a classical theorem, which describes the eigenvalues and left eigenvectors of a multiplication map. We refer to [9, Théorème 4.23] for details.
@@
In the situation of Theorem 3.1, if $A$ has rank n1n-1 and $I_{A,b}=\langle g\rangle$, then
@@ 4 Kuramoto oscillators
${\rm minimize}\,\sum_{j=1}^{n}\Big{(}x_{j}\,{\rm arcsin}(x_{j})+\sqrt{1-x_{j}^{2}}\Big{)},\quad{\rm subject\leavevmode\nobreak\ to}\quad Ax=\omega/K\,\,\text{and}\,\,x\in(-1,1)^{n}.$ (19)
@@
If $\omega/K\in\Omega^{+}$, then $A^{t}\theta^{*}\in{\rm int}(P)\subset{\rm Row}(A)$. In particular, $\cos(a_{l}\cdot\theta^{*})>0$ for all $l=1,\dotsc,n$. The statement then follows from [11, Lemma 3.2].
@@ 5 Positive points and convex optimization
$\dot{\theta}\,=\,-A\,\phi_{A,b}(\theta)+\omega,$ (20)
@@
The Lissajous curve ${\cal L}_{A,b}$ for the data $A=\begin{pmatrix}1&1\end{pmatrix}$, $b=(0,1)$ is the circle with defining equation $x^{2}+y^{2}=1$, see Example 1.2. The semi-algebraic set ${\cal L}_{A,b}^{+}$ is the segment of the curve contained in the (+,)(+,-) quadrant, see Figure 4.
@@
$\dot{\theta}\,=\,-\cos(\theta)-\sin(\theta)+\omega,\quad\omega\in\Omega_{A,b}^{+}$
has two steady state solutions $\theta^{*}$. Only one of them is stable. For concreteness, set $\omega=0.6$. The stable equilibrium is $\theta^{*}\approx-0.34725$. Its image $(\cos(\theta^{*}),\sin(\theta^{*}))$ is ${\cal L}_{A,b}^{+}\cap\{x+y=0.6\}$.
@@ 6 Lissajous discriminants
For $v={\bf 1}\in(\mathbb{C}^{*})^{d}$ and $\beta=-i\cdot{\bf 1}\in(\mathbb{C}^{*})^{n}$, we have $\det J_{A,{\bf 1}}({\bf 1})=\det(-i\,A\,A^{t})\neq 0$ because $A$ has full rank. This shows the first statement. For the second part of the lemma, since $\beta$ is generic we may assume that $\det J_{A,b}(v)$ is not identically zero as a Laurent polynomial in $v$. Pick $v_{0}\in(\mathbb{C}^{*})^{d}$ such that $\det J_{A,b}(v_{0})\neq 0$ and let $\omega_{0}=A\psi_{A,b}(v_{0})$. By construction, $v_{0}$ is isolated in ${\rm pr}_{\omega}^{-1}(\omega_{0})$. Since ${\rm pr}_{\omega}$ is a dominant morphism of irreducible $d$-dimensional varieties, this implies that its generic fiber is finite [17, Chapter 1, §8, Theorem 2 and Corollary 1].
@@
Here we start from an ideal with $d+1$ generators in $\mathbb{C}[v^{\pm 1},\omega]$. We compute two examples.
@@
$\Delta_{A,\bf 0}=$ $-8\,\omega_{1}^{5}+4\,\omega_{1}^{4}\,\omega_{2}^{2}-20\,\omega_{1}^{4}\,\omega_{2}-23\,\omega_{1}^{4}+8\,\omega_{1}^{3}\,\omega_{2}^{3}-8\,\omega_{1}^{3}\,\omega_{2}^{2}-46\,\omega_{1}^{3}\,\omega_{2}+4\,\omega_{1}^{3}+4\,\omega_{1}^{2}\,\omega_{2}^{4}$
@@
Different regions in Figure 6 correspond to different numbers of real solutions to (22), or different numbers of real intersection points in ${\cal L}_{A,b}\cap\{Ax=\omega\}$. For $\omega$ in the connected component of $\mathbb{R}^{2}\setminus\nabla_{A,\bf 0}$ highlighted in orange, the line $Ax=\omega$ intersects ${\cal C}_{A}$ in three real points. In all other connected components, there is only one real intersection point. On the right, the green, blue and red components correspond to $\omega$ with six, four and two real intersection points respectively. We note that the convex hull of the real points of $\nabla_{A,\bf 1}$ contains the blue region $\Omega^{+}$ in Figure 3 (right), but the two do not coincide.
@@
If the Lissajous discriminant variety $\nabla_{A,b}$ is a hypersurface, then it has degree at most $(\deg{\rm pr}_{\omega})^{-1}\cdot d\cdot d!\,{\rm vol}(P_{A})$. The number ${\rm vol}(P_{A})$ is as in Theorem 2.5.
@@
The degree of the Lissajous discriminant can be computed by solving (25) for generic $\omega_{0},\omega_{1}$. There are $\delta$ solutions, whose image under ${\rm pr}_{\omega}$ consists of $\deg\nabla_{A,b}$ many points. To compute all $\delta$ solutions, we use the Julia package HomotopyContinuation.jl (v2.15.0) [4]. To compute the lattice volume $d!\,{\rm vol}(P_{A})$, we use the function lattice_volume from Oscar.jl (v1.1.2) [18]. We apply these methods for matrices $A$ coming from cyclic and complete graphs. The results are reported in Table 1. The column $d\cdot d!\,{\rm vol}(P_{A})$ is computed using the command d*lattice_volume(convex_hull(transpose([A -A]))). By [5, Theorem 14], these numbers for $C_{n}$ are given by the sequence $n(n-1)\binom{n-1}{\lfloor\tfrac{1}{2}(n-1)\rfloor}$. The other entries can be reproduced using the following Julia code snippet, which uses standard tools:

### v2

@@ 1 Introduction
Let $A=\begin{pmatrix}a_{1}&a_{2}&\cdots&a_{n}\end{pmatrix}\in\mathbb{Z}^{d\times n}$ be such that ${\rm rank}(A)=d$ and $\mathbb{Z}A=\mathbb{Z}^{d}$. Let $P_{A}\subset\mathbb{R}^{d}$ be the polytope obtained as the convex hull of the lattice points $\{\pm a_{1},\ldots,\pm a_{n}\}\subset\mathbb{Z}^{d}$. For any $b\in\mathbb{C}^{n}$, the variety ${\cal L}_{A,b}$ is irreducible of dimension $d$. Moreover, for generic $b\in\mathbb{C}^{n}$, the degree of ${\cal L}_{A,b}$ is $\deg({\cal L}_{A,b})=2^{-{\rm CL}_{A}}\,d!\,{\rm vol}(P_{A})$, where ${\rm vol}(\cdot)$ denotes the euclidean volume and ${\rm CL}_{A}$ is the number of zero entries of a generic vector in the kernel of $A:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$.
The quantity $d!\,{\rm vol}(P_{A})$ is called the normalized volume of $P_{A}$. The polytope $P_{A}$ in Example 1.2 is the line segment [1,1][-1,1], with normalized volume two, which is the degree of the circle. The polygon $P_{A}$ in Example 1.3 is a hexagon with normalized volume $6=\deg{\cal S}_{A}$. The notion of “generic bb” in Theorem 1.4 will be made more precise in Section 2, and we give a formula for $\deg({\cal L}_{A,b})$ which holds for any $A,b$, but requires more notation (Theorem 2.5). We show that the hypersurface ${\cal C}_{A}$ obtained from the incidence matrix $A$ of the $n$-cycle graph $C_{n}$ is the cycle polynomial from [25, Section 4.2]. Using Theorem 2.5 and a result from [5], we prove a formula for the degree of the cycle polynomial (Proposition 2.10).
@@ 2 Dimension and degree
It remains to show that $S_{v}$ consists of $d!\,{\rm vol}(P_{A})$ points. This is an application of Kushnirenko’s theorem [13, Théorème III’], which predicts the maximal (and expected) number of solutions to a system of Laurent polynomial equations with identical monomial support. In order to apply this theorem, we must ensure that our equations are non-degenerate with respect to the polytope $P_{A}$. This means that none of the facial subsystems have a solution in $(\mathbb{C}^{*})^{d}$. By our assumption that ${\rm rank}(A)=d$, the centrally symmetric polytope $P_{A}$ is full-dimensional and it contains the origin in its interior. This implies that no face of $P_{A}$, except $P_{A}$ itself, contains both the lattice points $a_{j}$ and $a_{j}$. Therefore, there are no dependencies among the coefficients appearing in any facial subsystem. Hence, for generic $c_{jk}$, our equations are indeed non-degenerate and Kushnirenko’s upper bound $|S_{v}|\leq d!\,{\rm vol}(P_{A})$ is attained.
@@
We end the section by applying our degree formula to a family of Chebyshev hypersurfaces which arises when studying elliptopes of cycle graphs [23, 25]. This is relevant in semidefinite completion problems, see for instance [15]. Rephrasing [25, Equation (28)] in our notation, the $n$-th cycle polynomial $\Gamma_{n}^{\prime}$ is the defining equation of the Chebyshev hypersurface ${\cal C}_{A_{n}}$, with
@@
Notice that $\Gamma_{n}^{\prime}$ is defined up to scale. Its degree was computed up to $n=11$ in [25, Table 1]. In the paragraph preceding Conjecture 4.9 in [25], it is stated that a closed formula for $\deg(\Gamma_{n}^{\prime})$ is not known. The next proposition provides such a formula.
@@ 3 Determinantal equations
Fix $A\in\mathbb{Z}^{d\times n}$ and $b\in\mathbb{C}^{n}$. Let $\beta\in(\mathbb{C}^{*})^{n}$ as usual, and $Y_{A,\beta}\subseteq\mathbb{C}^{n}$ be the corresponding scaled affine toric variety. If $g_{1},\dotsc,g_{r}$ generate the ideal of $Y_{A,\beta}\cap(\mathbb{C}^{*})^{n}$ in $\mathbb{C}[y_{1}^{\pm 1},\ldots,y_{n}^{\pm 1}]$, then
@@
It is easier to obtain generators for the ideal of $Y_{A,\beta}\cap(\mathbb{C}^{*})^{n}$ than for the toric ideal of $Y_{A,\beta}$. First, one computes a $\mathbb{Z}$-kernel of $A$. This can be found, for instance, via the Smith normal form of $A$. If this kernel is generated by $m_{1},\ldots,m_{r}\in\mathbb{Z}^{n}$, with $r=n-{\rm rank}(A)$, then the binomials $g_{1},\ldots,g_{r}$ in Theorem 3.1 are given by $y^{m_{k}}-\beta^{m_{k}},k=1,\ldots,r$ [26]. To avoid computing matrix inverses, one may clear denominators by writing $m_{k}=u_{k}-w_{k}$ with $u_{k},w_{k}\in\mathbb{N}^{n}$ and use $g_{k}=\beta^{w_{k}}y^{u_{k}}-\beta^{u_{k}}y^{w_{k}}$ instead.
@@
To prove Theorem 3.1, we first state a classical theorem, which describes the eigenvalues and left eigenvectors of a multiplication map. We refer to [10, Théorème 4.23] for details.
@@
In the situation of Theorem 3.1, if $A$ has rank n1n-1 and $I_{A,\beta}=\langle g\rangle$, then
@@ 4 Kuramoto oscillators
${\rm minimize}\,\sum_{j=1}^{n}\Big(x_{j}\,{\rm arcsin}(x_{j})+\sqrt{1-x_{j}^{2}}\Big),\quad{\rm subject\penalty 10000\ to}\quad Ax=\omega/K\,\,\text{and}\,\,x\in(-1,1)^{n}.$ (19)
@@
If $\omega/K\in\Omega^{+}$, then $A^{t}\theta^{*}\in{\rm int}(P)\subset{\rm Row}(A)$. In particular, $\cos(a_{l}\cdot\theta^{*})>0$ for all $l=1,\dotsc,n$. The statement then follows from [12, Lemma 3.2].
While solving the equations (17) leads to all potential equilibria, there may be many “spurious” solutions which have no physical meaning. Proposition 4.7 says that if $\omega/K\in\Omega^{+}$, then a linearly stable solution can be found more efficiently via convex optimization. We now consider a larger graph to illustrate the computational range of this technique.
Example 4.9.
The graph shown in Figure 4 appears in [12, Example 3.1]. The corresponding Lissajous variety ${\cal S}_{A}$ has degree 832. For $\omega=(0.1,\ldots,0.1)$, the unique minimizer in (19) is
$x^{*}\,=\,(-0.050,0.150,0.083,0.017,-0.050,0.150,0.017,0.083,-0.067,0.200,0.067,0.200),$
here rounded to three decimal places. This can be computed using, for instance, JuMP.jl [17]. The optimum is found in less than 0.01 seconds. Solving $A^{t}\theta^{*}={\rm arcsin}(x^{*})$ we find that
$\theta^{*}=(0.151,0.285,0.151,0.285,0.201,0.268,0.201),$
which is a linearly stable equilibrium of the Kuramoto equations. We verify that $x\in{\cal S}_{A}^{+}\cap\{A\,x=\omega\}$ and $\omega\in\Omega^{+}$ by checking that $\omega=A\,{\rm sin}(A^{t}\theta^{*})$, and $A^{t}\theta^{*}\in[-\tfrac{\pi}{2},\tfrac{\pi}{2}]^{12}$.
Figure 4: A graph and its reduced incidence matrix.
@@ 5 Positive points and convex optimization
$\dot{\theta}\,=\,A\,\phi_{A,b}(\theta)-\omega,$ (20)
@@
The Lissajous curve ${\cal L}_{A,b}$ for the data $A=\begin{pmatrix}1&1\end{pmatrix}$, $b=(0,1)$ is the circle with defining equation $x^{2}+y^{2}=1$, see Example 1.2. The semi-algebraic set ${\cal L}_{A,b}^{+}$ is the segment of the curve contained in the (,+)(-,+) quadrant, see Figure 5.
@@
$\dot{\theta}\,=\,\cos(\theta)+\sin(\theta)-\omega,\quad\omega\in\Omega_{A,b}^{+}$
has two steady state solutions $\theta^{*}$. Only one of them is stable. For concreteness, set $\omega=0.6$. The stable equilibrium is $\theta^{*}\approx 1.918$. Its image $(\cos(\theta^{*}),\sin(\theta^{*}))$ is ${\cal L}_{A,b}^{+}\cap\{x+y=0.6\}$.
@@ 6 Lissajous discriminants
For $v={\bf 1}\in(\mathbb{C}^{*})^{d}$ and $\beta=-i\cdot{\bf 1}\in(\mathbb{C}^{*})^{n}$, we have $\det J_{A,{\bf 1}}({\bf 1})=\det(-i\,A\,A^{t})\neq 0$ because $A$ has full rank. This shows the first statement. For the second part of the lemma, since $\beta$ is generic we may assume that $\det J_{A,b}(v)$ is not identically zero as a Laurent polynomial in $v$. Pick $v_{0}\in(\mathbb{C}^{*})^{d}$ such that $\det J_{A,b}(v_{0})\neq 0$ and let $\omega_{0}=A\psi_{A,b}(v_{0})$. By construction, $v_{0}$ is isolated in ${\rm pr}_{\omega}^{-1}(\omega_{0})$. Since ${\rm pr}_{\omega}$ is a dominant morphism of irreducible $d$-dimensional varieties, this implies that its generic fiber is finite [19, Chapter 1, §8, Theorem 2 and Corollary 1].
@@
Here we start from an ideal with $d+1$ generators in $\mathbb{C}[v^{\pm 1},\omega]$. We compute three examples.
@@
[…]
@@
Different regions in Figure 7 correspond to different numbers of real solutions to (22), or different numbers of real intersection points in ${\cal L}_{A,b}\cap\{Ax=\omega\}$. For $\omega$ in the connected component of $\mathbb{R}^{2}\setminus\nabla_{A,\bf 0}$ highlighted in orange, the line $Ax=\omega$ intersects ${\cal C}_{A}$ in three real points. In all other connected components, there is only one real intersection point. On the right, the green, blue and red components correspond to $\omega$ with six, four and two real intersection points respectively. We note that the convex hull of the real points of $\nabla_{A,\bf 1}$ contains the blue region $\Omega^{+}$ in Figure 3 (right), but the two do not coincide. Figure 8 shows a comparison. This is similar to the inclusion $(-1,1)\subsetneq(-\sqrt{2},\sqrt{2})$ in Example 6.7.
@@
Figure 8: Comparing $\Omega^{+}$ (blue) with the Lissajous discriminant (purple).
@@
If the Lissajous discriminant variety $\nabla_{A,b}$ is a hypersurface, then it has degree at most $(\deg{\rm pr}_{\omega})^{-1}\cdot d\cdot d!\,{\rm vol}(P_{A})$, where ${\rm pr}_{\omega}$ denotes the restriction $({\rm pr}_{\omega})_{|R_{A,b}}$. The number ${\rm vol}(P_{A})$ is as in Theorem 2.5.
@@
The degree of the Lissajous discriminant can be computed by solving (25) for generic $\omega_{0},\omega_{1}$. There are $\delta$ solutions, whose image under ${\rm pr}_{\omega}$ consists of $\deg\nabla_{A,b}$ many points. To compute all $\delta$ solutions, we use the package HomotopyContinuation.jl (v2.15.0) [4] in Julia (v1.11). To compute the lattice volume $d!\,{\rm vol}(P_{A})$, we use the function lattice_volume from Oscar.jl (v1.1.2) [8, 20]. We apply these methods for matrices $A$ coming from cyclic and complete graphs. The results are reported in Table 1. The column $d\cdot d!\,{\rm vol}(P_{A})$ is computed using the command d*lattice_volume(convex_hull(transpose([A -A]))). By [5, Theorem 14], these numbers for $C_{n}$ are given by the sequence $n(n-1)\binom{n-1}{\lfloor\tfrac{1}{2}(n-1)\rfloor}$. The other entries can be reproduced using the following Julia code snippet, which uses standard tools: